Part 1) The critical value of t at a .01 level in a one-tail test will equal the critical value of t for a two-tail test at what level of significance?

  1. .05
  2. .10
  3. .20
  4. None of these

    Part 2) The critical value of t at a .05 level in a two-tail test will equal the critical value of t for a one-tail test at what level of significance?

  5. .025
  6. .10
  7. .20
  8. None of these
Answer #1

Part 1

Answer: d. None of these

For the given condition, critical value for α = 0.01 for one-tailed test is equal to the critical value for α = 2*0.01 = 0.02 for two-tailed test. So, answer is 0.02. Answer is not listed in the alternatives.

Part 2

Answer: e. 0.025

For the given condition, critical value for α = 0.05 for two-tailed test is equal to the critical value for α = 0.05/2 = 0.025 for one-tailed test. So, answer is 0.025.